Two cyclists start at the same point and travel in opposite directions. One cyclist travels 8 mph slower than the other. If the two cyclists are 136 miles apart after 4 hours, what is the rate of each cyclist?
What is the rate of the slower cyclist?
What is the rate of the faster cyclist?

Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(53765)

.

Let x be the rate of the slower cyclist, in miles per hour.

Then the rate of the faster cyclist is (x+8) mph.


The total distance equation in 4 hours is


    4x + 4*(x+8) = 136   miles apart


Simplify and solve.


    4x + 4x + 4*8 = 136

    8x            = 136 - 32

    8x            = 104

     x            = 104/8 = 13 mph.


ANSWER.  The slower cyclist rate is 13 miles per hour.

Solved.

---------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.



Answer by Boreal(15235)

Their combined speed apart is 136/4 =34 mph
x and x+8 are the speeds
2x+8=34
2x=26
x=13 mph slower cyclist
x+8=21 mph faster cyclist
in 4 hours, they go 52 and 84 miles respectively.